OpenStudy (anonymous):

2x+1=-4+7x showing all work

OpenStudy (anonymous):

\[2x+1=-4+7x\] \[-5x=-5\] x=1

OpenStudy (anonymous):

2x+1=-4+7x 2x to the other side of the equals sign leaves you with: 1=-4+5x (you've subtracted 2x now) then you move the -4 to the other side, which leaves you with: -1-4=5x, which is: -5=5x -5/5=x